List the maximum and minimum for the following system of equations.
x is greater than or equal to 0
y is greater than or equal to 0
y is smaller than or equal to x+10
y is smaller than or equal to 2x+1
f(x)=x-2y

Answer by robertb(5830)

The feasibility region is an unbounded region with the following corner points:
(0,0), (1, 19), and (0,1).
By the fundamental theorem of linear programming, the max and min are among these corner points.
At (0,0), f = 0-2*0 = 0
At (1, 19), f = 1 - 2*-19 = -37
At (0,1), f = 0 - 2*1 = -2
Therefore the maximum value is 0 and the minimum value is -37.